# Making the OST of the Muon Collider

### v1

@@ I Introduction
Since the pioneering deep inelastic scattering (DIS) experiments at SLAC, the repertoire of phenomena that can be used as tools to teach us about the partonic structure of visible matter has been steadily expanded by improved accelerator facilities. Lepton scattering at high energies and with high integrated luminosity provides the best opportunity to understand the partonic structure of visible matter. Processes with large cross sections have already been measured, yet our understanding of nucleon structure in terms of partons remains largely one-dimensional. The future of hadronic physics therefore lies in measuring processes with small cross sections but unique sensitivity to otherwise inaccessible physical phenomenaĀ AchenbachĀ etĀ al. (2024).
@@
Historically, energy and luminosity have been anti-correlated for lepton-scattering experiments. For example, the collider experiments H1 and ZEUS at HERA operated at GeV but were limited to around 1 fb-1 of integrated luminosity accumulated over 15 years of running, while fixed-target experiments at much lower energies such as those at Jefferson Lab ( GeV) regularly collect a similar integrated luminosity in seconds. This is attributable mainly to the fact that even dilute gaseous targets are many orders of magnitude denser than collider beams.
@@
Multi-dimensional study of hadrons necessitates high luminosity to combat the curse of dimensionality. At the same time, high energy is vital to reach the regime where perturbative QCD is valid and predictive, enabling extraction of nucleon structure quantities from perturbative calculations. The instantaneous luminosity of the EIC is planned to be to cm-2/s, which is sufficient for multi-dimensional measurements of processes with low cross sections that were inaccessible at HERAĀ AbdulĀ KhalekĀ etĀ al. (2022). However, this luminosity is still significantly lower than those accessible at fixed-target experiments. Additionally, the present scope of the EIC includes a single collider experiment. Since the capabilities of the EIC machine are unique, there will be no ability to independently cross check the results of that experiment unless a second detector is constructed.
@@
With these aspects in mind, one can ask what other facilities can complement or surpass the EIC in terms of energy and luminosity. Perhaps the most similar existing facility to the EIC in terms of energy and luminosity is the M2 beamline of the CERN SPS, home to AMBER and COMPASSĀ AdamsĀ etĀ al. (2018); AbbonĀ etĀ al. (2007). M2 provides a secondary beam of muons with energies up to 190 GeV ( GeV) and attained a luminosity of around cm-2/s for COMPASS. The highest energy fixed-target lepton-scattering experiment to date was E665 at Fermilab, which operated with a muon beam of around 500 GeVĀ AdamsĀ etĀ al. (1990) but a luminosity below cm-2/s. Two future proposals located at CERN, the LHeCĀ AgostiniĀ etĀ al. (2021); BordryĀ etĀ al. (2018) and FCC-ehĀ BordryĀ etĀ al. (2018); AbadaĀ etĀ al. (2019), reach higher energies by adding electron rings to the LHC and FCC. Another future facility that has been discussed is a Muon-Ion ColliderĀ AcostaĀ andĀ Li (2022); ClineĀ E. (2022); AcostaĀ etĀ al. (2025); DavoudiaslĀ etĀ al. (2024) that could reach a of between 140 GeV and 6.5 TeV with a luminosity of to cm-2/s, lower than the EIC and similar to HERA.
@@
The goal of this paper is to present another future alternative to complement and expand upon the physics of the EIC in the mid-21st century using the beam of the Muon Collider for a fixed-target experiment. We denote this experiment Muons On Stationary Targets, or OST. The Muon Collider concept has been discussed for decades, culminating in positive endorsements for Muon Collider R&D by the most recent iterations of the European Particle Physics StrategyĀ CER (2020) group and U.S. Particle Physics Project Prioritization PanelĀ AsaiĀ etĀ al. (2023). The 2025 U.S. National Academies report on elementary particle physicsĀ National Academies of Sciences, Engineering, and Medicine (2025) listed Muon Collider R&D as Recommendation 1 for U.S. particle physics in the coming decades. As the recently formed International Muon Collider Collaboration undertakes R&D and design studies, it is timely to consider the inherent potentiality of the Muon Collider for experiments in nuclear and neutrino physics111The idea of using the Muon Collider beam as a source of neutrinos for fixed-target experiments has been discussed widely; a specific example is the nuSTORM conceptĀ AdeyĀ etĀ al. (2015). The high-intensity beams of TeV-scale neutrinos would also be a powerful tool for hadronic structure measurements..
@@ II Energy and Luminosity
Various center-of-mass energies have been explored as possible operation points for Muon Colliders, ranging from the Higgs pole at 62.5 GeV per beam to 7 TeV per beam. For simplicity, we restrict ourselves to considering the designs discussed in Ref.Ā AccetturaĀ etĀ al. (2023). The 7 TeV beam provides the highest energy and lowest beam-induced background for a fixed-target experiment; however, it makes the design of the detector more challenging due to the highly boosted center-of-mass, as will be described in Sec.Ā III. The physics reach can be extended by performing measurements also at lower beam energies, where the final state is less forward-boosted.
@@
Projections for the beam conditions of the Muon Collider were provided in Ref.Ā AccetturaĀ etĀ al. (2023), and we will use these numbers to make some rough estimates of what could be achieved in terms of luminosity in a fixed-target experiment. For the 7 TeV beam, the number of muons per bunch is and the repetition rate is 5 Hz, corresponding to an integrated beam current of around 1.4 . This is about two orders of magnitude smaller than the beam current available at Jefferson Lab, but for muon scattering experiments, the target can be made very long without significantly impacting the muons as they enter and leave the target. If this beam were delivered directly onto a 1-meter-long liquid hydrogen target, it would produce approximately 3.3 ab-1 of integrated luminosity per day, resulting in an average instantaneous luminosity of 4 cm-2/s.
@@
While the aforementioned luminosity scenario appears attractive at first, the high bunch charge and low repetition rate will lead to enormous pileup if the structure of the beam reaching the target is the same as it is in the collider. One bunch of muons at 7 TeV passing through a 1 m long liquid hydrogen target could result in tens of thousands of simultaneous scattering events reaching the detector system. Realistically, the beam will only be dumped once a significant fraction of the muons have decayed. However, even if only 1% of the muons sent into the collider are put on target, the potential physics case is still significant, as will be shown in Sec.Ā IV.
@@
We assume this ā1% scenarioā for several reasons. The extraction of 1% of the collider beam is unlikely to affect the statistical precision attainable by the collider experiments. With 1% of the nominal collider beam current and an appropriately chosen detector acceptance, the level of pileup is similar to the HL-LHC. With the assumption of additional R&D advancements in detector and data processing technologies driven by HL-LHC, FCC, and the Muon Collider detectors, this seems like a technically feasible operating point. The pileup issue can be further ameliorated by extracting muons for the fixed-target experiment every turn of the collider instead of dumping the beam all at once, but this adds complexity to the collider ring design. Assuming 20 weeks per year of physics running, the 1% beam dump on a 1-meter target scenario would result in an average instantaneous luminosity of approximately 1.5 cm-2/s, around an order of magnitude higher than the maximum luminosity of the EIC. Operation at beam energies down to 1.5 TeV should provide similar luminosities for a fixed-target experiment, due to the similar repetition rate and number of muons per bunchĀ 222In the collider, the increase in luminosity with energy is driven by the smaller beam size and number of useful rotations, neither of which significantly affect the fixed-target luminosity.. Furthermore, depending on the dump design, the and beams could be utilized independently by two experiments located at the and beam dumps, respectively. A longer target, such as the 6-meter-long target of EMCĀ AllkoferĀ etĀ al. (1981), could potentially increase the luminosity to almost 1 cm-2/s, but at the cost of higher pileup.
@@
The authors of Ref.Ā CesarottiĀ etĀ al. (2023) demonstrated the impressive reach of a beam dump experiment for studying long-lived BSM particles. Since the muons are highly penetrating, such a beam dump experiment could be operated in parallel with the nuclear physics experiment described here. With the assumed 1% scenario, the long-lived particle search experiment would reach muons on target in the course of a couple years of running.
@@ III Detector, Target, and Environment
Several aspects of the unique muon beam environment have significant impact on the design of the detector systems. One can imagine the detector taking several forms. A large-aperture forward detector such as LHCb or COMPASS is an obvious baseline. We assume this experiment could not be built in the collider ring Ć la LHCb or HERMES due to the immense beam-induced background and the space required to precisely momentum analyze the high-energy scattered muons, which scatter at small angles even at high momentum transfer, . Therefore, we assume that the detector resides in a long hall with a dedicated beamline in the manner of COMPASS or SeaQuest. A rough sketch of such a detector with particle identification (PID) via Cherenkov is shown in Fig.Ā 2.
@@
Perhaps the most obvious challenge is that at 7 TeV the lab-frame angles of produced particles are extremely small even at high . Therefore it seems necessary to utilize a dipole magnet that can appreciably bend particles at small scattering angles. The momentum resolution of a detector with a dipole analyzing magnet can be estimated as , where is the spatial resolution of the detector, is the distance from the magnet at which the measurement is performed, is the magnetic field strength, and is the distance traveled in the magnetic field. To have acceptable resolution on the DIS variables within a reasonable range in inelasticity , the momentum and angle of the scattered muon should be measured precisely even at TeV energies. LHCb achieves momentum resolution of around 1% for 200 GeV particles with a of 4 Tm; to achieve a similar momentum resolution for 2 TeV particles with the same detector would require a of 40 Tm. Perhaps a more practical solution is to increase the lever arm, , compared to LHCb, which is around 20 m long. For a longer spectrometer, it could be feasible to have a dipole magnet of smaller angular aperture than LHCb with of 20 Tm and a factor of two larger lever arm. The dipole could be made superconducting to reach higher field, at the cost of increased engineering complexity. This momentum resolution coupled with a scattering angle resolution of 2%, similar to what has been achieved in COMPASSĀ AdamsĀ etĀ al. (2018), would enable measurements of at the 4% level in the region GeV2 for the 7 TeV beam.
Another option, perhaps better suited for inclusive and semi-inclusive DIS measurements, is to utilize relatively small-acceptance magnetic spectrometers such as those used in Halls A and C at Jefferson LabĀ BeneschĀ etĀ al. (2022); AliĀ etĀ al. (2025) and in End Station A at SLAC. The physically smaller magnets in these spectrometers can more easily reach higher -fields than large aperture magnets, providing substantially better momentum resolution for the scattered muon or high momentum hadrons. The drawback of this option, aside from the reduced acceptance, is that pileup can result in large numbers of random in-time coincidences for coincidence experiments unless the vertex location can be precisely determined. Augmentation of the classical spectrometer design with precise vertex detectors is likely necessary, but since the muons are less affected by multiple scattering, this should not significantly impact the spectrometer resolution. While this small-acceptance spectrometer setup may be a more cost-effective option, we will focus on the large-acceptance detector design.
@@
The beam-induced background (BIB) plays a dominant role in the design of the detector for Muon Collider experiments. In the collider, this background arises from secondary particles produced by decay electrons interacting with the beamline and surrounding materials. For 7 TeV muons at the nominal collider beam current of per bunch, there will be approximately decays per meterĀ AccetturaĀ etĀ al. (2023), rising to at 1.5 TeV. The current plan for the collider experiments is to install tungsten cones near the beamline that absorb secondary particles and use fast timing detectors to eliminate out-of-time BIB. For collider experiments the luminosity is typically inversely proportional to the distance between the set of accelerator magnets after the interaction region, meaning longer detectors reduce the luminosity. There is no such constraint on fixed-target experiments, and experimental halls are typically on the order of 20-50 m long. This means that the detector can be shielded from BIB produced upstream of the target with several meters of shielding. Depending on the distance to the experimental hall from the collider ring, it could be preferable to have a final bending magnet immediately upstream of the experiment that bends the primary beam into the experimental hall and dumps the decay electrons, which carry approximately one-third of the beam momentum on average. The BIB that may be challenging to completely mitigate is the decay electrons produced inside and after the target. These electrons can be deflected out of the beamline and into the detector acceptance by the dipole magnet. For this reason, it is likely necessary to magnetically shield the primary beamline from the detector dipole as was done in HERMESĀ AckerstaffĀ etĀ al. (1998), at the cost of experimental acceptance.
@@
Vertexing, fast timing, and advanced data acquisition techniques such as online reconstruction for triggering will be crucial to mitigate pileup and BIB, particularly for longer and wider target options. Fast-timing detectors will be able to reject hits from out-of-time BIB and reduce the occupancy; significant R&D and experience in this direction can be inherited from future collider experiments. For heavy-flavor measurements, the produced bottom or charm mesons will likely decay inside the target, meaning the vertex detector should ideally be able to distinguish displaced vertices inside the target. The LHCb collaboration has demonstrated software-based triggering capable of implementing complex analysis cuts at the trigger-levelĀ AaijĀ etĀ al. (2019). Since scattering events are kinematically over-constrained when the hadronic final state is measured, this kind of streaming data acquisition pipeline will be an extremely powerful tool to maximize the amount of useful data collected in a harsh environment.
One optional configuration for the detector system is to employ a technique often used in hadron beam Drell-Yan experiments, such as SeaQuest, whereby thick absorbers filter out all final-state particles except for muons. This could be a more cost-effective day-one configuration of the detector with a full LHCb-style detector being implemented later on. In the detector design with an absorber, scattered muons and muons from hadron decays would have significantly lower backgrounds. Triple-muon events could be used to study highly sought-after processes such as double-DVCSĀ GuidalĀ andĀ Vanderhaeghen (2003) and electroproduction of vector mesons, e.g. and , for imaging the spatial distributions of partons inside of nucleons and nuclei.

### v2

# A Fixed-Target Experiment at the Muon Collider

@@ I Introduction
Since the pioneering deep inelastic scattering (DIS) experiments at SLAC, the repertoire of phenomena that can be used as tools to teach us about the partonic structure of visible matter has been steadily expanded by improved accelerator facilities. Lepton scattering at high energies and with high integrated luminosity provides the best opportunity to understand the partonic structure of visible matter. Processes with large cross sections have already been measured, yet our understanding of nucleon structure in terms of partons remains largely one-dimensional. The future of hadronic physics therefore lies in measuring processes with small cross sections but unique sensitivity to otherwise inaccessible physical phenomena Achenbach:2023pba; AbdulKhalek:2021gbh.
@@
Historically, energy and luminosity have been anti-correlated for lepton-scattering experiments (see Fig. 1). For example, the collider experiments H1 and ZEUS at HERA operated at GeV but were limited to around 1 fb-1 of integrated luminosity accumulated over 15 years of running, while fixed-target experiments at much lower energies such as those at Jefferson Lab ( GeV) regularly collect a similar integrated luminosity in seconds. This is attributable mainly to the fact that even dilute gaseous targets are many orders of magnitude denser than collider beams.
@@
Multi-dimensional study of hadrons necessitates high luminosity to combat the curse of dimensionality. At the same time, high energy is vital to reach the regime where perturbative QCD is valid and predictive, enabling extraction of nucleon structure quantities from perturbative calculations. The instantaneous luminosity of the EIC is planned to be to cm-2/s, which is sufficient for multi-dimensional measurements of processes with low cross sections that were inaccessible at HERA AbdulKhalek:2021gbh. However, this luminosity is still significantly lower than those accessible at fixed-target experiments. Additionally, the present scope of the EIC includes a single collider experiment. Since the capabilities of the EIC machine are unique, there will be no ability to independently cross check the results of that experiment unless a second detector is constructed.
@@
With these aspects in mind, one can ask what other facilities can complement or surpass the EIC in terms of energy and luminosity. Perhaps the most similar existing facility to the EIC in terms of energy and luminosity is the M2 beamline of the CERN SPS, home to AMBER and COMPASS Adams:2018pwt; COMPASS:2007rjf. M2 provides a secondary beam of muons with energies up to 190 GeV ( GeV) and attained a luminosity of around cm-2/s for COMPASS. The highest energy fixed-target charged-lepton-scattering experiment to date was E665 at Fermilab, which operated with a muon beam of around 500 GeV E665:1989lkq but a luminosity below cm-2/s. Two future proposals, the LHeC LHeC:2020van; Bordry:2018gri and FCC-eh Bordry:2018gri; FCC:2018byv, reach higher energies by adding electron rings to the LHC and FCC. Another future facility that has been discussed is a Muon-Ion Collider Acosta:2021qpx; Acosta:2022ejc; ClineMuIC; Acosta:2025tdh; Davoudiasl:2024fiz; Apyan:2025ujv; Zhao:2025kvq; Ahluwalia:2022qsp that could reach a of 140 GeV to 6.5 TeV with a luminosity of to cm-2/s, lower than the EIC and similar to HERA.
@@
The goal of this paper is to present another future alternative to complement and expand upon the physics of the EIC in the mid-21st century using the beam of the Muon Collider for a fixed-target experiment. We denote this experiment Muons On Stationary Targets, or OST. The Muon Collider concept has been discussed for decades Ruggiero:1992zv; Proceedings:1993ere; Barger:1995hr; NeutrinoFactory:2005cgg; Mokhov:2011zzd; Delahaye:2013jla; Wang:2015yyh; MuonCollider:2022nsa, culminating in positive endorsements for Muon Collider R&D by the most recent iterations of the European Particle Physics Strategy CERN-ESU-015-2020 group and U.S. Particle Physics Project Prioritization Panel P5:2023wyd. The 2025 U.S. National Academies report on elementary particle physics NAS2025Higgs listed Muon Collider R&D as Recommendation 1 for U.S. particle physics in the coming decades. As the recently formed International Muon Collider Collaboration InternationalMuonCollider:2025sys undertakes R&D and design studies, it is timely to consider the inherent potentiality of the Muon Collider for hadronic physics111The idea of using the Muon Collider beam as a source of neutrinos for fixed-target experiments has been discussed widely; a specific example is the nuSTORM concept Adey:2015iha. High-intensity beams of TeV-scale neutrinos are also inherently powerful for hadronic structure measurements NuTeV:2001whx; Bai:2020ukz; Anchordoqui:2021ghd; Feng:2022inv; Xie:2023suk; Dulat:2015mca; Candido:2023utz; deGouvea:2025zfq; Ariga:2025qup; InternationalMuonCollider:2025sys; Francener:2025pnr..
@@ II Energy and Luminosity
Various center-of-mass energies have been explored as possible operation points for Muon Colliders, ranging from the Higgs pole at 62.5 GeV per beam to 7 TeV per beam. For simplicity, we restrict ourselves to considering the designs discussed in Ref. Accettura:2023ked. The 7 TeV beam provides the highest energy and lowest beam-induced background for a fixed-target experiment; however, it makes the design of the detector more challenging due to the highly boosted center-of-mass, as will be described in Sec. III. The physics reach can be extended by performing measurements also at lower beam energies, where the final state is less forward-boosted.
@@
Projections for the beam conditions of the Muon Collider were provided in Ref. Accettura:2023ked, and we will use these numbers to make some rough estimates of what could be achieved in terms of luminosity in a fixed-target experiment. For the 7 TeV beam, the number of muons per bunch is and the repetition rate is 5 Hz, corresponding to an integrated beam current of around 1.4 . This is about two orders of magnitude smaller than the beam current available at Jefferson Lab, but for muon scattering experiments, the target can be made very long without significantly impacting the muons as they enter and leave the target. If this beam were delivered directly onto a 1-meter-long liquid hydrogen target, it would produce approximately 3.3 ab-1 of integrated luminosity per day, resulting in an average instantaneous luminosity of 4 cm-2/s.
@@
While the aforementioned luminosity scenario appears attractive at first, the high bunch charge and low repetition rate will lead to enormous pileup if the structure of the beam reaching the target is the same as it is in the collider. One bunch of muons at 7 TeV passing through a 1 m long liquid hydrogen target could result in tens of thousands of simultaneous scattering events reaching the detector system. Realistically, the beam will only be dumped once a significant fraction of the muons have decayed. However, even if only 1% of the muons sent into the collider are put on target, the potential physics case is still significant, as will be shown in Sec. IV.
@@
We assume this “1% scenario” for several reasons. The extraction of 1% of the collider beam is unlikely to affect the statistical precision attainable by the collider experiments. With 1% of the nominal collider beam current and an appropriately chosen detector acceptance, the level of pileup is similar to the HL-LHC. With the assumption of additional R&D advancements in detector and data processing technologies driven by HL-LHC, FCC, and the Muon Collider detectors, this seems like a technically feasible operating point. The pileup issue can be further ameliorated by extracting muons for the fixed-target experiment every turn of the collider instead of dumping the beam all at once, but this adds complexity to the collider ring design. Assuming 20 weeks per year of physics running, the 1% beam dump on a 1-meter target scenario would result in an average instantaneous luminosity of approximately 1.5 cm-2/s, around an order of magnitude higher than the maximum luminosity of the EIC. Operation at beam energies down to 1.5 TeV should provide similar luminosities for a fixed-target experiment, due to the similar repetition rate and number of muons per bunch 222In the collider, the increase in luminosity with energy is driven by the smaller beam size and number of useful rotations, neither of which significantly affect the fixed-target luminosity.. Furthermore, depending on the dump design, the and beams could be utilized independently by two experiments located at the and beam dumps, respectively. A longer target, such as the 6-meter-long target of EMC EuropeanMuon:1980nje333A more extreme example is the 40-meter target of BCDMS Bollini:1982ac., could potentially increase the luminosity to almost 1 cm-2/s, but at the cost of higher pileup.
@@
The authors of Refs. Cesarotti:2022ttv; Cesarotti:2023sje demonstrated the impressive reach of a beam dump experiment for studying long-lived BSM particles. Since the muons are highly penetrating, such a beam dump experiment could be operated in parallel with the nuclear physics experiment described here. With the assumed 1% scenario, the long-lived particle search experiment would reach muons on target in the course of a couple years of running.
@@ III Detector, Target, and Environment
Several aspects of the unique muon beam environment have significant impact on the design of the detector systems. One can imagine the detector taking several forms. A large-aperture forward detector such as LHCb or COMPASS is an obvious baseline. We assume this experiment could not be built in the collider ring à la LHCb or HERMES due to the immense beam-induced background and the space required to precisely momentum analyze the high-energy scattered muons, which scatter at small angles even at high momentum transfer, . Therefore, we assume that the detector resides in a long hall with a dedicated beamline in the manner of COMPASS COMPASS:2007rjf or SeaQuest SeaQuest:2017kjt. A rough sketch of such a detector with particle identification (PID) via Cherenkov is shown in Fig. 2.
@@
Perhaps the most obvious challenge is that at 7 TeV the lab-frame angles of produced particles are extremely small even at high . Therefore it seems necessary to utilize a dipole magnet that can appreciably bend particles at small scattering angles. The momentum resolution of a detector with a dipole analyzing magnet can be estimated as , where is the spatial resolution of the detector, is the distance from the magnet at which the measurement is performed, is the magnetic field strength, and is the distance traveled in the magnetic field. To have acceptable resolution on the DIS variables within a reasonable range in inelasticity , the momentum and angle of the scattered muon should be measured precisely even at TeV energies. LHCb achieves momentum resolution of around 1% for 200 GeV particles with a of 4 Tm; to achieve a similar momentum resolution for 2 TeV particles with the same detector would require a of 40 Tm. Perhaps a more practical solution is to increase the lever arm, , compared to LHCb, which is around 20 m long. For a longer spectrometer, it could be feasible to have a dipole magnet of smaller angular aperture than LHCb with of 20 Tm and a factor of two larger lever arm. The dipole could be made superconducting to reach higher field, at the cost of increased engineering complexity. This momentum resolution coupled with a scattering angle resolution of 2%, similar to what has been achieved in COMPASS Adams:2018pwt, would enable measurements of at the 4% level in the region GeV2 for the 7 TeV beam.
@@
The beam-induced background (BIB) plays a dominant role in the design of the detector for Muon Collider experiments. In the collider, this background arises from secondary particles produced by decay electrons interacting with the beamline and surrounding materials. For 7 TeV muons at the nominal collider beam current of per bunch, there will be approximately decays per meter Accettura:2023ked, rising to at 1.5 TeV. Fixed-target experimental halls are typically on the order of 20-50 m long. This means that the detector can be shielded from BIB produced upstream of the target with several meters of shielding. Depending on the distance to the experimental hall from the collider ring, it could be preferable to have a final bending magnet immediately upstream of the experiment that bends the primary beam into the experimental hall and dumps the decay electrons, which carry approximately one-third of the beam momentum on average. The BIB that may be challenging to completely mitigate is the decay electrons produced inside and after the target. These electrons can be deflected out of the beamline and into the detector acceptance by the dipole magnet. For this reason, it is likely necessary to magnetically shield the primary beamline from the detector dipole as was done in HERMES HERMES:1998mat, at the cost of experimental acceptance.
@@
Vertexing, fast timing, and advanced data acquisition techniques such as online reconstruction for triggering will be crucial to mitigate pileup and BIB, particularly for longer and wider target options. Fast-timing detectors will be able to reject hits from out-of-time BIB and reduce the occupancy; significant R&D and experience in this direction can be inherited from future collider experiments MuonCollider:2022glg; MAIA:2025hzm; InternationalMuonCollider:2025sys. For heavy-flavor measurements, the produced bottom or charm mesons will likely decay inside the target, meaning the vertex detector should ideally be able to distinguish displaced vertices inside the target. The LHCb collaboration has demonstrated software-based triggering capable of implementing complex analysis cuts at the trigger-level LHCb:2018zdd. Since scattering events are kinematically over-constrained when the hadronic final state is measured, this kind of streaming data acquisition pipeline will be an extremely powerful tool to maximize the amount of useful data collected in a harsh environment.
One optional configuration for the detector system is to employ a technique often used in hadron beam Drell-Yan experiments, such as SeaQuest SeaQuest:2017kjt, whereby thick absorbers filter out all final-state particles except for muons. This could be a more cost-effective day-one configuration of the detector with a full LHCb-style detector being implemented later on. In the detector design with an absorber, scattered muons and muons from hadron decays would have significantly lower backgrounds. Triple-muon events could be used to study highly sought-after processes such as double-DVCS Guidal:2002kt and electroproduction of vector mesons, e.g. and , for imaging the spatial distributions of partons inside of nucleons and nuclei.